OpenStudy (anonymous):

This is a problem on proportions: 2 bugs start from zero on a number line and they move at constant speeds in opposite directions. After 4 sec P reaches 12 and N reaches -8. a)If x is the point reached by P after t sec and y is the point reached by N after t seconds, express x and y in terms of t.

OpenStudy (anonymous):

Please help!

OpenStudy (phi):

)If x is the point reached by P after t sec write a proportion \[ \frac{x}{t} = \frac{12}{4} \] you can simplify that to \[ \frac{x}{t} =3 \] if you multiply both sides by t, and simplify you get \[ x = 3t \]

OpenStudy (anonymous):

Thank you so very much!

OpenStudy (anonymous):

Also will y be -2t or +2t ? Thanks

OpenStudy (phi):

y = -2t as a test, they tell us y is the point reached by N after t seconds, and after 4 sec N reaches -8 using 4 for t in the equation y= -2 t we get y = -2*4 = -8 which matches what they told us.

OpenStudy (anonymous):

Thanks again!